OpenStudy (anonymous):

4/x =y/5, what is the value of y/x? A) 1/20 B) 4/5 C) 5/4 D) 20 E) cannot be determine from the information given.

OpenStudy (saifoo.khan):

Insert the pairs.. it will let u knw

myininaya (myininaya):

\[\frac{4}{x}=\frac{y}{5}=> 4*5=y*x=>20=xy=> \frac{20}{x}=\frac{y}{x}\] not enough info

myininaya (myininaya):

we have enough info for yx or xy? but not enough for y/x

OpenStudy (anonymous):

ohh okay ty :)